OpenStudy (elleblythe):

Solve for x: Arctan(1-x)-Arctanx=pi

OpenStudy (anonymous):

\[\tan^{-1}(a) - \tan^{-1}(b) = \tan^{-1}({\frac{a - b}{1 + ab})}\]

OpenStudy (perl):

wolfram says no solutions exist http://www.wolframalpha.com/input/?i=Arctan%281-x%29-Arctanx%3Dpi

OpenStudy (anonymous):

tan pi=0, so, 1-2x=0, so x=0.5

OpenStudy (anonymous):

x= 1/2 should not be the solution.. :)

OpenStudy (anonymous):

Put in back in the equation, \(0 \ne \pi\) ??

OpenStudy (anonymous):

arctan((1-2x)/(1+x-x^2))=pi, so it should. Where isn't it satisfying again?

OpenStudy (perl):